a t.v is marked 25% above the cost price.in order to gain 10% what should be the percentage of discount allowed on it?
Answer by fcabanski(1391)

This will be easier if you pick a number for c. Pick a simple number c=100.


If the cost was 100, then a 25% (.25) profit comes from 125. That's 100(1+.25).


10% profit comes from 110.
125(1+x)=110 where x is the discount number we seek.


125 + 125x = 110


125x=-15


x = -15/125 = -.12. We need a 12% discount on the $125 TV in order to obtain $110, which is a $10 profit. 125*(1-.12) = 110